# LSMS: Language-guided Scale-aware MedSegmentor for Medical Image Referring Segmentation

### v2

@@ Abstract
Conventional medical image segmentation methods have been found inadequate in facilitating physicians with the identification of specific lesions for diagnosis and treatment. Given the utility of text as an instructional format, we introduce a novel task termed Medical Image Referring Segmentation (MIRS), which requires segmenting specified lesions in images based on the given language expressions. Due to the varying object scales in medical images, MIRS demands robust vision-language modeling and comprehensive multi-scale interaction for precise localization and segmentation under linguistic guidance. However, existing medical image segmentation methods fall short in meeting these demands, resulting in insufficient segmentation accuracy. In response, we propose an approach named Language-guided Scale-aware MedSegmentor (LSMS), incorporating two appealing designs: (1) a Scale-aware Vision-Language Attention module that leverages diverse convolutional kernels to acquire rich visual knowledge and interact closely with linguistic features, thereby enhancing lesion localization capability; (2) a Full-Scale Decoder that globally models multi-modal features across various scales, capturing complementary information between scales to accurately outline lesion boundaries. Addressing the lack of suitable datasets for MIRS, we constructed a vision-language medical dataset called Reference Hepatic Lesion Segmentation (RefHL-Seg). This dataset comprises 2,283 abdominal CT slices from 231 cases, with corresponding textual annotations and segmentation masks for various liver lesions in images. We validated the performance of LSMS for MIRS and conventional medical image segmentation tasks across various datasets. Our LSMS consistently outperforms on all datasets with lower computational costs. The code and datasets will be released.
1 Introduction
@@
Figure 1: Comparison between Medical Image Referring Segmentation and conventional medical image segmentation tasks, illustrating the distinction in feature handling process between the proposed LSMS and existing medical segmentation methods. In the images displaying segmentation results, the regions highlighted in red represent the Ground Truth. For intuitive correspondence with the left-right references in the text, all CT images in this paper have been mirrored horizontally.
@@
The significance of organ segmentation and lesion segmentation in medical image analysis has been widely recognized [29, 43]. As shown in Figure 1(a), the classical medical image segmentation task involves inputting a medical image and obtaining segmentation results encompassing all lesions within the image [3, 33]. With the advancement of multi-modal research, studies have emerged that incorporates textual input as supplementary information for segmentation. Figure 1(b) illustrates medical image segmentation with textual prompts, wherein diagnostic texts provided by physicians serve as language prompting. The task involves inputting a medical image along with its corresponding diagnostic text and outputting segmentation results for all lesions present [20]. However, in clinical practice, physicians often need to segment specific lesions to aid in diagnosis and treatment, thus rendering conventional medical image segmentation methods inadequate for practical applications. Language, as a convenient medium for expressing physicians’ needs, can be used to indicate target segmentation objects. Therefore, we introduce the task of Medical Image Referring Segmentation (MIRS), where medical images are accompanied by language expressions that indicate specific lesions within the images. Figure 1(c) represents the introduced task, MIRS, which aims to specifically output segmentation results for the lesion referenced in the language expressions. In the example depicted in the figure, segmentation of the largest lesion in the right liver is required.
Deep learning methods have demonstrated outstanding performance in image segmentation tasks. As illustrated in Figure 1(a), classical medical image segmentation often relies on architectures such as U-Net [29] or Transformer [32]. These approaches involve down-sampling feature maps to acquire high-level contextual information, followed by up-sampling to reconstruct spatial dimensions [3, 2]. To further enhance segmentation performance, approaches incorporating language prompting have emerged [20], as depicted in Figure 1(b). These methods fuse visual and linguistic features during the encoding stage of the model, aiding in the acquisition of additional contextual information. The latest referring segmentation methods in natural images also utilize the similar architecture [39]. However, due to the varying size and shape inherent in medical images, the fusion of single-scale representations at each stage with linguistic features alone is insufficient for precise object localization. Moreover, the sequential structure adopted in the decoding stage lacks a comprehensive understanding of features at various scales, thereby constraining segmentation performance.
@@
Upon analyzing the requirements of the MIRS task and previous efforts in medical image segmentation, we identify two primary challenges in MIRS: (i) Robust vision-language modeling. The considerable differences in size and shape among objects in the medical visual environment present a challenge for effectively integrating visual and linguistic features. Fusion between linguistic features and single-scale visual features may overlook the rich local visual information correlated with linguistic guidance, thereby impacting the model’s object localization performance. (ii) Comprehensive multi-scale interaction. Globally modeling the complex differences across scales enables the extraction of valuable global visual-linguistic features. Given the complexity of the visual environment in medical images, a lack of complementary information between scales may result in insufficient capability to identify lesion boundaries during segmentation.
@@
To address these challenges, we propose a novel MIRS architecture (feature handling process shown in Figure 1(c)), namely Language-guided Scale-aware MedSegmentor (LSMS). Within LSMS, we introduce a Scale-aware Vision-Language Attention (SVLA) module embedded in the encoder block. SVLA enhances the acquisition of scale-aware visual knowledge by employing convolutional kernels of various sizes, facilitating close interaction with linguistic features and thereby improving visual-linguistic consistency. As shown in Figure 2, LSMS (w/o FSD) equipped with SVLA exhibits a significant enhancement in lesion localization capability compared to the results of existing vision-language model LViT [20] and LAVT [39]. Additionally, we devise a Full-Scale Decoder (FSD) that globally models multi-modal information by aligning and integrating multi-modal feature maps from various scales, thereby enhancing the comprehension of details within complex medical visual environment.
@@
In Figure 2, LSMS, when contrasted with LSMS (w/o FSD), exhibits a more precise prediction of lesion boundaries during segmentation.
The proposed MIRS task lacks an available dataset, necessitating a corresponding language expression for the localization of each lesion in medical images. Therefore, we developed the Reference Hepatic Lesion Segmentation (RefHL-Seg) dataset, comprising 2,283 abdominal CT slices from 231 cases. With the assistance of radiologists, we annotated abdominal CT images containing liver lesions, providing language expressions for each lesion regarding its morphology, location, type, and other relevant information. Additionally, corresponding segmentation masks were delineated to facilitate training and validation for the MIRS task.
@@
    We introduce the novel task of MIRS, which entails locating and segmenting target objects in medical images based on reference expressions. The application of MIRS in clinical practice can enhance diagnostic efficiency for physicians.
@@
    We propose a model named LSMS for the MIRS task, comprising a SVLA module to improve object localization capability and a full-scale decoder to enhance the accuracy of lesion boundary prediction during segmentation.
@@
    We develop the Reference Hepatic Lesion Segmentation (RefHL-Seg) dataset for training and validating the MIRS task. RefHL-Seg consists of 2,283 abdominal CT slices from 231 cases.
@@
    We conduct comprehensive experiments on the RefHL-Seg dataset for MIRS, as well as on datasets for conventional medical image segmentation tasks. Experimental results demonstrate the superiority of LSMS over current state-of-the-art methods with lower computational costs.
Figure 3: An illustration of LSMS. Initially, the input image and the reference expression are embedded separately through the visual embedding block and the BERT [6] language encoder, yielding visual feature $V_{1}$ and linguistic feature $L$, which are then fed into the language-guided vision encoder. The encoder incorporates the Scale-aware Vision-Language Attention (SVLA) module to interact between visual knowledge from different receptive fields and linguistic features. The encoders at each stage learn rich multi-modal features $F_{i},i\in\{1,2,3,4\}$. Through the Position Alignment module, $F_{i},i\in\{1,2,3,4\}$ uniformly resize the feature maps of various scales while preserving channel disparities, resulting in $S_{i},i\in\{1,2,3,4\}$. Features $S_{i},i\in\{2,3,4\}$ are then globally modeled across scales by the full-scale decoder for final segmentation.
2 Related Works
2.1 Medical Image Segmentation
@@
For medical image segmentation tasks, early methods [22, 4, 38] extended from image classification networks to achieve semantic segmentation, among which the Fully Convolutional Network (FCN) [22] is an end-to-end semantic segmentation network [18]. Multi-scale architectures have been proven to enhance segmentation performance based on CNN methods. U-Net [29] is considered a pioneer in medical image segmentation. U-Net++ [43] further improved U-Net by enhancing skip connections to bridge the semantic gap between shallow encoder layers and deep decoder layers. MS-DualGuided [30] focused on both spatial and channel dimensions of feature maps at different scales. CA-Net [34] proposed a multi-scale context-aware network for multi-modal medical image segmentation. Subsequently, the introduction of Transformer [32] was found to enhance medical image segmentation performance. TransUNet [3] integrated the power of Transformers with the U-Net architecture for improved instance segmentation. Swin-Unet [2] specifically designed to leverage the hierarchical features and shifted window mechanisms of Swin Transformer [21], thereby enhancing the model’s capability to capture both local and global context. Furthermore, LViT [20] proposed a medical image segmentation model with language prompting. These methods acquire feature maps through progressively down-sampling, yielding low-resolution representations. And they utilize a sequential up-sampling structure in the decoding stage, lacking a global understanding of multi-scale knowledge.
2.2 Vision-Language Segmentation Models
The capability of vision-language models to encode rich multi-modal representations has played a crucial role in the field of computer vision [35, 42]. ViLT [17], by simplifying the processing of visual inputs in CLIP [28], proposed a more parameter-efficient architecture. This architecture enables interaction layers to handle visual features effectively without relying on separate deep visual embeddings. Subsequently, there have been significant research efforts in image segmentation [19], where text information is utilized to enhance model segmentation capabilities. VLT [7] and EFN [9] utilize a Transformer-based encoder-decoder framework, employing attention mechanisms in decoding stages to augment contextual information for natural image referring segmentation. LAVT [39] adopted an early fusion approach, modeling multi-modal context within the Vision Transformer (ViT) [8] encoder. Some methods [1, 24] have started using text information to assist in medical image analysis. [27] leveraged the knowledge transfer ability of pretrained vision-language models [18] to the medical domain by employing well-designed medical prompts. LViT [20] introduced textual prompts as auxiliary information during the encoding stage for segmentation. Due to the complexity of medical visual environment, existing methods struggle to achieve explicit alignment between text and images. To address this, we propose a design involving close interaction between scale-aware visual knowledge and linguistic features to facilitate modeling of visual-linguistic relationships.
3 Language-Guided Scale-Aware Medical Segmentor
3.1 Overview
@@
In the proposed Language-guided Scale-aware MedSegmentor (LSMS), we learn visual knowledge from diverse receptive fields and tightly engage with linguistic features. Subsequently, leveraging a full-scale decoder, we globally model multi-modal information across all scales, thereby facilitating MIRS task. The overall architecture of LSMS is presented in Figure 3.
@@
Given an image and a language expression, our model predicts the segmentation mask corresponding to the language reference. Our LSMS follows a workflow composed of [language-guided vision encoder] - [position alignment] - [full-scale decoder]. LSMS comprises four stages, each with varying numbers of encoder blocks and different feature map resolutions. The encoder (Section 3.2) incorporates a novel scale-aware vision-language attention module (Section 3.3), which learns rich visual knowledge through convolutions of different sizes and closely interacts with linguistic features. Additionally, we propose a full-scale decoder (Section 3.4) to globally model multi-modal information across multiple scales, enhancing the understanding of context details. In the following subsections, we elaborate on each component of LSMS.
3.2 Language-Guided Vision Encoder
@@
To facilitate deep interaction between linguistic features and the complex visual environment of medical images, our encoder leverages a novel attention module (Section 3.3) to perceive visual details in the image under linguistic guidance, thus obtaining valuable multi-modal representations.
Figure 4: (a) An illustration of the encoder block in the Language-Guided Vision Encoder. (b) An illustration of the Scale-aware Vision-Language Attention.
@@
As illustrated in Figure 3, our encoding phase comprises four stages with decreasing feature map resolutions. The $i$-th stage consists of $N_{i}$ encoder blocks. For the sake of clarity, we assume that $N_{i}=1,i\in\{1,2,3,4\}$ in this section. LSMS receives an input image alongside a reference expression. We extract the linguistic feature $L\in\mathbb{R}^{C_{l}\times T}$ using the language encoder BERT [6], where $T$ denotes the number of words, and $C_{l}$ represents the channel number of the linguistic feature. Similarly, the input image undergoes an embedding block to yield the initial visual input $V_{1}\in\mathbb{R}^{C_{v1}\times H_{1}\times W_{1}}$ for the encoder, where $H_{1}$ and $W_{1}$ represent the height and width of the visual feature, and $C_{v1}$ represents the channel number. Each stage comprises a down-sampling block and a stack of encoder blocks. For each stage, the aggregation of the multi-modal feature maps $F_{i}\in\mathbb{R}^{C_{vi}\times H_{i}\times W_{i}}$ can be expressed as:
$F_{i}=\begin{cases}LGVE(V_{1},L),&i=1\\ LGVE(Down(F_{i-1}),L),&i=2,3,4\\ \end{cases}$ (1)
where $i$ denotes the index of the stage, the function $Down(\cdot)$ represents the down-sampling block, and $LGVE(\cdot)$ denotes the encoder block. The visual input provided to the encoder block is obtained by $V_{i}=Down(F_{i-1})$. The down-sampling block consists of a convolutional layer with a stride of 2 and a kernel size of 3 × 3, followed by batch normalization.
@@
As depicted in Figure 4(a), the architecture of encoder blocks follows the design of ViT [8], but we introduce a SVLA module (Section 3.3) to replace the conventional self-attention mechanism. The workflow of the encoder block is illustrated by the following:
$F_{i}^{{}^{\prime}}=Norm(SVLA(V_{i},L)+V_{i}),$ (2)
$F_{i}=Norm(FeedForward(F_{i}^{{}^{\prime}})+F_{i}^{{}^{\prime}}),$ (3)
where $Norm(\cdot)$ and $FeedForward(\cdot)$ represent normalization and feed forward layers, $SVLA(\cdot)$ denotes SVLA module, and the output of $SVLA(V_{i},L)$ is labeled as $F_{i}^{Att}$.
3.3 Scale-aware Vision-Language Attention

### v3

# Language-guided Scale-aware MedSegmentor for Lesion Segmentation in Medical Imaging

@@ Abstract
In clinical practice, segmenting specific lesions based on the needs of physicians can significantly enhance diagnostic accuracy and treatment efficiency. However, conventional lesion segmentation models lack the flexibility to distinguish lesions according to specific requirements. Given the practical advantages of using text as guidance, we propose a novel model, Language-guided Scale-aware MedSegmentor (LSMS), which segments target lesions in medical images based on given textual expressions. We define this as a new task termed Referring Lesion Segmentation (RLS). To address the lack of suitable benchmarks for RLS, we construct a vision-language medical dataset named Reference Hepatic Lesion Segmentation (RefHL-Seg). LSMS incorporates two key designs: (i) Scale-Aware Vision-Language attention module, which performs visual feature extraction and vision-language alignment in parallel. By leveraging diverse convolutional kernels, this module acquires rich visual representations and interacts closely with linguistic features, thereby enhancing the model’s capacity for precise object localization. (ii) Full-Scale Decoder, which globally models multi-modal features across multiple scales and captures complementary information between them to accurately delineate lesion boundaries. Additionally, we design a specialized loss function comprising both segmentation loss and vision-language contrastive loss to better optimize cross-modal learning. We validate the performance of LSMS on RLS as well as on conventional lesion segmentation tasks across multiple datasets. Our LSMS consistently achieves superior performance with significantly lower computational cost. Code and datasets will be released.
Index Terms:
vision-language, medical image, segmentation, multi-scale
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
@@
Figure 1: Comparison between Referring Lesion Segmentation and conventional lesion segmentation tasks. In the images displaying segmentation results, the regions highlighted in red represent the Ground Truth. For intuitive correspondence with the left-right references in the text, all CT images in this paper have been mirrored horizontally.
@@
The significance of lesion segmentation in medical image analysis has been widely recognized [1, 2]. It enables the precise identification and delineation of pathological regions, which is essential for accurate diagnosis and treatment planning. As shown in Fig. 1(a), the Classical Lesion Segmentation task involves inputting a medical image and obtaining segmentation results encompassing lesions within the image [3, 4]. With the advancement of multi-modal research, studies have emerged that incorporates textual input as supplementary information for segmentation. Fig. 1(b) illustrates Text-Augmented Lesion Segmentation, wherein diagnostic texts provided by physicians aid in image interpretation [5]. This task involves inputting a medical image along with its diagnostic text and outputting segmentation results for All Lesions without distinction. However, in clinical practice, physicians often need to segment specific lesions to provide tailored diagnosis and treatment, thus rendering Conventional Lesion Segmentation  [^1] In this paper, Conventional Lesion Segmentation is used as a collective term to refer to the two tasks illustrated in Figure 1: (a) Classical Lesion Segmentation and (b) Text-Augmented Lesion Segmentation.
tasks inadequate for practical applications. For instance, in radiation therapy, the accuracy of tumor segmentation directly influences the precision of radiation beam targeting, which determines the effectiveness of the treatment and the potential risk of side effects. As shown in Fig. 1(a) and (b), Conventional Lesion Segmentation tasks treat all lesions equally, which is insufficient for clinical scenarios. Text, as a convenient medium for expressing physicians’ needs, can be used to indicate target segmentation objects. Therefore, we introduce a new task of Referring Lesion Segmentation (RLS), where medical images are accompanied by a language expression that indicate a Specific Lesion within the image. Fig. 1(c) represents the introduced task, RLS, which involves segmenting the largest lesion in the right liver as described by the language expression. Compared to the task in Fig. 1(b), RLS emphasizes the ability to differentiate lesions based on the text. To support evaluation of RLS, we developed Reference Hepatic Lesion Segmentation (RefHL-Seg) dataset—the first benchmark specifically designed for this new task.
Figure 2: Comparison of Transformer-based architectures. (a) Existing architectures for related tasks. (b) our LSMS for RLS.
Deep learning methods have demonstrated outstanding performance in medical image segmentation tasks. Classical Lesion Segmentation methods often relies on architectures such as U-Net [1] or Transformer [6]. Methods combining these two architectures generally down-sample feature maps to acquire high-level contextual information, followed by up-sampling to reconstruct spatial dimensions [3, 7]. To further enhance segmentation performance, approaches incorporating language modality have emerged [5, 8]. Recently, Transformer-based models have shown great promise in this area, as their ability to model long-range dependencies facilitates effective integration of visual and linguistic information. As shown in Fig. 2(a), existing Transformer-based methods include a fusion module at the end of each stage to integrate the extracted visual features with linguistic information. Models utilizing this architecture include the mature Natural Image Referring Segmentation model LAVT [9] and the recent Text-Augmented Lesion Segmentation model LViT [5]. By introducing a multi-scale structure, they effectively exploit rich vision-language knowledge. However, these methods are not directly applicable to the RLS task. Existing methods treat visual feature extraction and cross-modal fusion as two independent steps, leaving room for improvement in achieving visual-linguistic alignment within the semantic space. During decoding, they adopt sequential structures that tend to produce single-scale representations at each level, while enhancing inter-scale interaction may help capture core semantic information.
@@
Upon analyzing the requirements of RLS and previous efforts in related tasks, we identify two key aspects that address the unique challenges of RLS while aligning with conventional segmentation goals: (i) Robust vision-language modeling. In the medical visual environment, the notable variations in size and shape among objects make it crucial to effectively model visual-linguistic consistency for accurate object localization. Fusing visual and linguistic features after visual feature extraction may overlook the rich local visual information correlated with linguistic guidance, thereby impacting the model’s object localization performance. (ii) Comprehensive multi-scale interaction. Globally modeling the complex differences across scales enables the extraction of optimal global visual-linguistic features. Given the complexity of the visual environment in medical images, neglecting complementary information between scales may result in insufficient capability to identify lesion boundaries during segmentation.
@@
In light of the aforementioned analysis, we propose a model named Language-guided Scale-aware MedSegmentor (LSMS), as illustrated in Fig. 2(b). Within LSMS, we introduce a Scale-aware Vision-Language Attention (SVLA) module embedded in the encoder block. SVLA captures scale-aware visual knowledge and models visual-linguistic relationships in an integrated manner, enhancing the visual-linguistic alignment in the semantic space. As shown in Fig. 3, LSMS (w/o FSD) equipped with SVLA exhibits a significant enhancement in lesion localization capability compared to the results of existing vision-language model LViT [5] and LAVT [9]. Additionally, we devise a Full-Scale Decoder (FSD) that globally models multi-modal information by aligning and integrating multi-modal feature maps from various scales, thereby enhancing the comprehension of details within complex medical images. In Fig. 3(b), LSMS, when contrasted with LSMS (w/o FSD), exhibits a more precise prediction of lesion boundaries during segmentation.
@@
Additionally, we have designed a specialized loss function to constrain the model’s training, which includes the Segmentation Loss $\mathcal{L}_{Seg}$ for the segmentation results and the Vision-Language Contrastive Loss $\mathcal{L}_{Con}$ for the linguistic features and final multi-modal features. Specifically, $\mathcal{L}_{Seg}$ enhances the focus on the object boundaries, thereby improving edge detection capabilities. Meanwhile, $\mathcal{L}_{Con}$ further aligns the visual knowledge of the target region with the linguistic information, leading to improved accuracy in target localization.
@@
    We introduce a new task of RLS, which entails segmenting the target object in medical images based on the language expression. We have established the RefHL-Seg dataset as a new benchmark for RLS.
@@
    We propose LSMS for lesion segmentation, comprising a SVLA module to improve object localization capability and a full-scale decoder to enhance the accuracy of lesion boundary prediction.
@@
    We design a specialized loss function to optimize fine-grained discrimination and visual-linguistic alignment, thereby enhancing the accuracy.
@@
    We conduct comprehensive experiments on the RefHL-Seg dataset for RLS, as well as on datasets for conventional lesion segmentation tasks. Experimental results demonstrate the superiority of LSMS over current state-of-the-art methods with lower computational costs.
The preliminary work was presented as a conference paper at the International Joint Conference on Artificial Intelligence (IJCAI) 2023 [10]. This paper extends the prior work by introducing a new task (RLS), a dataset (RefHL-Seg) as the new benchmark, as well as method and experimental enhancements. We optimize the scale-wise module design in the model and propose a specialized loss function, which improves the model’s performance in the domain of lesion segmentation. Furthermore, we apply the proposed model to the newly introduced RLS task and validate it on conventional lesion segmentation benchmarks, demonstrating the generalizability of our approach.
II Related Works
II-A Medical Image Segmentation
@@
For medical image segmentation tasks, early methods [11, 12] extended from image classification networks to achieve semantic segmentation, among which the Fully Convolutional Network (FCN) [11] is an end-to-end semantic segmentation network [18]. Multi-scale architectures have been proven to enhance segmentation performance based on CNN methods. U-Net [1] is considered a pioneer in medical image segmentation. U-Net++ [2] further improved U-Net by enhancing skip connections to bridge the semantic gap between shallow encoder layers and deep decoder layers. MS-DualGuided [13] focused on both spatial and channel dimensions of feature maps at different scales. CA-Net [14] proposed a multi-scale context-aware network for multi-modal medical image segmentation. Subsequently, the introduction of Transformer [6] was found to enhance medical image segmentation performance. TransUNet [3] integrated the power of Transformers with the U-Net architecture for improved instance segmentation. Swin-Unet [7] specifically designed to leverage the hierarchical features and shifted window mechanisms of Swin Transformer [15], thereby enhancing the model’s capability to capture both local and global context. Furthermore, LViT [5] proposed a medical image segmentation model with textual assistance, introducing text input as auxiliary information during the encoding stage of segmentation. RecLMIS [8] proposes a novel cross-modal conditioned method for Text-Augmented Lesion Segmentation. While current research incorporates text as an aid for image understanding, existing benchmarks and methods are unable to distinguish specific objects in medical images, thereby falling short of fully addressing the practical needs of physicians.
Figure 4: (a) An illustration of LSMS. Initially, the input image and the reference expression are embedded separately through the vision embedding block and the BERT [16] language encoder, yielding visual feature $V_{1}$ and linguistic feature $L$, which are then fed into the language-guided vision encoder. The encoder incorporates the Scale-aware Vision-Language Attention (SVLA) module to interact between visual knowledge from different receptive fields and linguistic features. The encoder blocks at each stage learn rich multi-modal features $F_{i},i\in\{1,2,3,4\}$, which are subsequently fed into the full-scale decoder. Through the Position Alignment block, $F_{i},i\in\{1,2,3,4\}$ uniformly resize the feature maps of various scales while preserving channel disparities, resulting in $P_{i},i\in\{1,2,3,4\}$. Features $P_{i},i\in\{2,3,4\}$ are then globally modeled across scales for final segmentation. (b) An illustration of the operational mechanism of the vision-language contrastive loss $\mathcal{L}_{Con}$.
II-B Natural Image Referring Segmentation
Early referring segmentation methods in natural images [17, 18, 19] typically combined visual and linguistic features by concatenation, utilizing FCNs for cross-modal feature learning and prediction. In contrast, attention-based methods, such as vision-guided linguistic attention [20] and Cross-Modal Self-Attention [21], focus on aligning visual content with linguistic information. Bi-directional relationship networks [22] capture mutual guidance, while others [23, 24] use sentence structure to model cross-modal attributes. More recently, Transformer-based methods have enhanced the modeling of long-range cross-modal dependencies, leading to significant advancements in referring segmentation. VLT [25] and EFN [26] utilize a Transformer-based encoder-decoder framework, employing attention mechanisms in decoding stages to augment contextual information. LAVT [9] adopt an early fusion approach, modeling multi-modal context within the Transformer encoders. Prompt-RIS [27] leverages bidirectional prompting for better vision-language interaction. Current methods for Natural Image Referring Segmentation treat visual feature extraction and cross-modal fusion as two separate steps, and use sequential upsampling structures during decoding. However, due to the complexity of medical visual environments, these methods struggle to meet the requirements of RLS. In this paper, we propose the SVLA, which models visual information and visual-linguistic relationships in an integrated manner, and employ a full-scale decoder to promote comprehensive understanding of multi-scale knowledge.
III Language-Guided Scale-Aware Medical Segmentor
III-A Overview
@@
In the proposed Language-guided Scale-aware MedSegmentor (LSMS), we learn visual knowledge from diverse receptive fields and tightly engage with linguistic features. Subsequently, leveraging a full-scale decoder, we globally model multi-modal information across all scales, thereby facilitating RLS task. The training loss includes the Segmentation Loss $\mathcal{L}_{{Seg}}$ and the vision-language contrastive loss $\mathcal{L}_{{Con}}$. The overall architecture of LSMS is presented in Fig. 4.
@@
Given an image and a language expression, our model predicts the segmentation mask corresponding to the language reference. Our LSMS follows a workflow composed of [language-guided vision encoder] - [full-scale decoder]. LSMS comprises four stages, each with varying numbers of encoder blocks and different feature map resolutions. The encoder (Sec. III.B) incorporates a novel Scale-aware Vision-Language Attention module (Sec. III.C), which learns rich visual knowledge through convolutions of different sizes and closely interacts with linguistic features. Additionally, we propose a full-scale decoder (Sec. III.D) to globally model multi-modal information across multiple scales, enhancing the understanding of context details. During training, we employ the $\mathcal{L}_{{Seg}}$ to supervise the segmentation results and the $\mathcal{L}_{{Con}}$ to constrain the linguistic features and the final multi-modal representations (Sec. III.E). In the following subsections, we elaborate on each component of LSMS.
III-B Language-Guided Vision Encoder
@@
To facilitate deep interaction between linguistic features and the complex visual environment of medical images, our encoder leverages a novel attention module (Sec. III.C) to perceive visual details in the image under linguistic guidance, thus obtaining valuable multi-modal representations. The structure of the encoder block is depicted in Fig. 5(a).
@@
As illustrated in Fig. 4, our encoding phase comprises four stages with decreasing feature map resolutions. The $i$-th stage consists of $N_{i}$ encoder blocks. For the sake of clarity, we assume that $N_{i}=1,i\in\{1,2,3,4\}$ in this section. LSMS receives an input image alongside a reference expression. We extract the linguistic feature $L\in\mathbb{R}^{C_{l}\times T}$ using the BERT [16] language encoder, where $T$ denotes the number of words, and $C_{l}$ represents the channel number of the linguistic feature. Similarly, the input image undergoes an embedding block to yield the initial visual input $V_{1}\in\mathbb{R}^{C_{v1}\times H_{1}\times W_{1}}$ for the encoder, where $H_{1}$ and $W_{1}$ represent the height and width of the visual feature, and $C_{v1}$ represents the channel number. Each stage comprises a down-sampling block and a stack of encoder blocks. For each stage, the aggregation of the multi-modal feature maps $F_{i}\in\mathbb{R}^{C_{vi}\times H_{i}\times W_{i}}$ can be expressed as:
[…]
@@
As depicted in Fig. 5(a), the architecture of encoder blocks follows the design of ViT [28], but we introduce a SVLA module (Sec. III.C) to replace the conventional self-attention mechanism. The workflow of the encoder block is illustrated by the following:
[…]
TABLE I: Detailed settings of different stages in our LSMS.
Stage 1 Stage 2 Stage 3 Stage 4
number of blocks ($N_{i}$) 3 3 5 2
channel 64 128 320 512
III-C Scale-aware Vision-Language Attention